Operations Report Dec 07th – 12th

MDRS Operations Report 7-12- Dec-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Electrician will be here either Monday or Tuesday

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 50.08

Minimum 47.10

Maximum 58.40

Notes on solar: We are adjusting power use to compensate for the weather. We had two days of overcast weather and SOC got to 16% but on sunny days and with our conservation efforts, we are still running on solar alone.

Propane Reading, station tank – 59 % %

Propane Reading, director tank— 66 %

Propane Reading, intern tank— 38%

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 400

Static to Loft Pump used – yes

Water in GreenHab – 250 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 500 gallons

Hab toilet tank emptied: yes

Perseverance rover used: yes

Hours: 225.4

Beginning charge: 100

Ending charge: 69

Currently charging: no

Sojourner rover used: yes

Hours: 158.1

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: Yes

Hours: 145.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: No

Hours: 75.6

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: Yes

Hours: 154.9

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Percy was rolled onto its driver’s side on Saturday at Robert’s Rock Garden. A tow company was called to upright it and a damaged rim was fixed enough for us to get the rover back to the station. Current assessed damage is the front driver’s tire rim, which will need to be replaced, and structural damage to the plastic outer parts. There may be some readjustment required for the bed. The batteries and “engine” appear fine, the brakes and gears are operational and it drives normally. Further assessment will be done in the next few days and any other damage will be reported in our next report. Percy has been removed from the station and is at the Outpost pending repairs.

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town to get water

CrewCar used and why, where? No. Battery will be replaced early this wek.

Luna used and why, where? Yes, around MDRS

General notes and comments: Nothing to report

Summary of internet: Crew internet was down for a day but has been restored. All other accounts nominal.

Summary of suits and radios: All unplugged except for communication radios

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Hab is fully operational with 4 people staying there. Both the furnace and the wall heater are being used for heat. This way we can conserve power overnight as the wall heater does not use power.

Summary of Outpost operations: Work continues on the Intern Trailer for a move in date of December 18th. Currently the water system is frozen as the wall heater was not on overnight. We will inspect for broken pipes as soon as the water thaws.

Summary of GreenHab operations: The new LED lights have been installed and we will be using supplemental lighting again beginning tonight from 7 pm to midnight.

Summary of ScienceDome operations: Used for science teaching during the recent three day visit with Ogden Prep.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Observatory was used for a short time during the class visit. There is currently a mount problem with the Robotic Observatory.

Summary of health and safety issues: Minor injuries reported during rover rollover.

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site